# New Applications of Ergodic Theory to Sets of Differences

### v1

@@
Abstract. We apply the methods of ergodic theory to both simplify and significantly extend some classical results due to Stewart, Tijdeman, and Ruzsa. One of the novel features of our approach is the utilization of pointwise ergodic theory.
@@ 1 Introduction
$\overline{d}(S)=\limsup_{N\rightarrow\infty}\frac{|S\cap\{1,\ldots,N\}|}{N},$ (1.1)
@@
$\underline{d}(S)=\liminf_{N\rightarrow\infty}\frac{|S\cap\{1,\ldots,N\}|}{N}.$ (1.2)
If $\overline{d}(S)=\underline{d}(S)$ then their common value is denoted by $d(S)$ and is called the natural density of $S$. Following Rusza we will introduce 3 types of difference sets; note that these difference sets are symmetric subsets of $\mathbb{Z}$.
@@
Suppose that $S_{1},\ldots,S_{k}\subseteq\mathbb{N}$ and $\overline{d}(S_{i})>0$ for $1\leq i\leq k$. In [17] Stewart and Tijdeman answered in the affirmative a question of Erdős by showing that the set $C\vcentcolon=\bigcap_{i=1}^{k}\Delta_{3}(S_{i})$ has positive lower density. Moreover they showed $C$ is syndetic, meaning that there are integers $m_{1},\ldots,m_{\ell}$ such that
@@
Stewart and Tijdeman’s result was amplified by Rusza as follows:
Theorem 1.1 ([15, Theorem 1 and Theorem 2]). if $S_{1},\ldots,S_{k}\subseteq\mathbb{N}$ satisfy $\overline{d}(S_{i})>0$ for each $i=1,\dots,k$, then there exists $S\subseteq\mathbb{N}$ such that $d(S)\geq\prod_{i=1}^{k}\overline{d}(S_{i})$ and $\Delta_{1}(S)\subseteq D\vcentcolon=\bigcap_{i=1}^{k}\Delta_{2}(S_{i})$. Furthermore, there are $m_{1},\ldots,m_{\ell}\in\mathbb{Z}$, where
@@
The goal of this paper is to offer an ergodic approach to Theorem 1 which not only provides an alternative proof, but also leads to various generalizations in the setup of countably infinite amenable cancellative [^2] A semigroup $G$ is (two-sided) cancellative if, for any $a,b,c\in G$, $ab=ac\Longrightarrow b=c$ and $ba=ca\Longrightarrow b=c$. semigroups. Our ergodic approach also leads to variants of Theorem 1 which involve modified versions of sets of differences $\Delta_{1}(S)$ and $\Delta_{2}(S)$ such as
$\Delta_{1,c}(S)=\{n:S\cap(S-[n^{c}])\neq\emptyset\}\quad\text{ and }\quad\Delta_{2,c}(S)=\{n:\overline{d}(S\cap(S-[n^{c}]))>0\}$ (1.5)
with $c>0$ and $c\notin\mathbb{N}$, or more generally, for $g:\mathbb{N}\to[0,\infty]$
$\Delta_{1,g}(S)=\{n:S\cap(S-[g(n)])\neq\emptyset\}\quad\text{ and }\quad\Delta_{2,g}(S)=\{n:\overline{d}(S\cap(S-[g(n)]))>0\}$ (1.6)
(In formulas (1.5) and (1.6), []delimited-[][$\cdot$][ ⋅ ] denotes the integral part.) For example, the following theorem is an immediate corollary of Theorem 3.2, to be proved in Subsection 3.2, which not only holds for $n^{c}$, but applies to a rather extensive family of functions coming from Hardy fields.
@@
Theorem 1.2. Let $c>0$, $c\not\in\mathbb{N}$. If $S_{1},\dots,S_{k}\subseteq\mathbb{N}$ satisfy $\overline{d}(S_{i})>0$ for each $i=1,\dots,k$, then there is $S\subseteq\mathbb{N}$ with $d(S)\geq\prod_{i=1}^{k}\overline{d}(S_{i})$ and $\Delta_{1,c}(S)\subseteq D_{c}\vcentcolon=\bigcap_{i=1}^{k}\Delta_{2,c}(S_{i})$. Furthermore, we have that $\overline{d}(D_{c})\geq\prod_{i=1}^{k}\overline{d}(S_{i})$.
@@
The following theorem is one of the results generalizing Theorem 1 which will be proved in this paper. We observe that Theorem 1 generalizes Theorem 1 in more than one way. First, Theorem 1 extends the set-up from $\mathbb{Z}$ to a general countably infinite amenable group $G$. Second, for the sets $S_{1},\dots S_{k}$, which appear in the formulation it is assumed $\overline{d}_{\mathbf{F}_{i}}(S_{i})>0$ for $i=1,\dots k$, where the Følner sequences $\mathbf{F}_{1},\dots,\mathbf{F}_{k}$ are arbitrarily chosen. Finally, the tempered Følner sequence $\mathbf{G}$, which appears in the formulation and whose precise definition is introduced and discussed in Subsection 4.2, leads even in the case $G=\mathbb{Z}$ to a nontrivial generalization of Theorem 1.
@@
Theorem 1.3. (Theorem 4.2, Subsection 4.2) Let $G$ be a countably infinite amenable group with (left) Følner sequences $\mathbf{G},\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$ where $\mathbf{G}$ is (left) tempered. Let $S_{1},\ldots,S_{k}\subseteq G$ such that $\overline{d}_{\mathbf{F}_{i}}(S_{i})>0$ for each $i=1,\ldots,k$. Then there is $S\subseteq G$ such that $d_{\mathbf{G}}(S)\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$ and $\Delta_{1}(S)\subseteq D=\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i})$. Furthermore, there are $m_{1},\ldots,m_{\ell}\in G$ where
@@
The structure of this paper is as follows. In Section 2 we provide a short ergodic proof of Theorem 1. The goal of Section 2 is also to set-up some ideas that will be further developed and amplified in subsequent sections. In Section 3 we obtain a general result involving the difference sets defined in (1.6) which contains Theorem 1 as an instantaneous corollary. In Section 4 we will prove Theorem 1 and its more general version pertaining to cancellative amenable semigroups. Finally, in Section 5 we give examples of tempered Følner sequences in $(\mathbb{N},\cdot)$, the Heisenberg group, and in locally finite groups.
@@ 2 Ergodic proof of Ruzsa’s Theorem
The proof of Theorem 1 we present in this section will utilize two facts from ergodic theory. The first of them is an extension of the classical Poincaré recurrence theorem  [^3] Poincaré’s recurrence theorem states that if $(X,\mathcal{B},\mu)$ is a probability space, $T$ a $\mathcal{B}$-measurable map $X\to X$ such that $\mu(T^{-1}A)=\mu(A)$ for all $A\in\mathcal{B}$, and $B\in\mathcal{B}$ with $\mu(B)>0$, then there is $n\in\mathbb{N}$ such that $\mu(B\cap T^{-n}B)>0$..
Theorem 2.1 ([3, Theorem 1.2]). Let $(X,\mathcal{B},\mu,T)$ be a measure-preserving system [^4] A measure-preserving system is a quadruple $(X,\mathcal{B},\mu,T)$ where $(X,\mathcal{B},\mu)$ is a probability space and $T$ is a $\mathcal{B}$-measurable map $X\to X$ such that for any $A\in\mathcal{B}$, one has $\mu(T^{-1}A)=\mu(A)$. A measure-preserving system $(X,\mathcal{B},\mu,T)$ is invertible if the map $T$ is invertible. and let $B\in\mathcal{B}$ with $\mu(B)>0$. Then there is a set $P\subseteq\mathbb{N}$ with $d(P)\geq\mu(B)$ such that for any $n_{1},\ldots,n_{r}\in P$ we have $\mu\left(B\cap T^{-n_{1}}B\cap\cdots\cap T^{-n_{r}}B\right)>0$.
@@
Theorem 2.3 (cf. [4, Theorem 1.1]). Suppose that $S\subseteq\mathbb{N}$ with $\overline{d}(S)>0$. Then there is an invertible measure-preserving system $(X,\mathcal{B},\mu,T)$ and a set $B\in\mathcal{B}$ with $\mu(B)=\overline{d}(S)$ such that
@@
Proof of Theorem 1. By Theorem 2, for each $i=1,\ldots,k$ there exists an invertible measure-preserving system $(X_{i},\mathcal{B}_{i},\mu_{i},T_{i})$ and a set $B_{i}\in\mathcal{B}_{i}$ with $\mu_{i}(B_{i})=\overline{d}(S_{i})$ such that $S_{i}$, $B_{i}$, and $T_{i}$ satisfy equation (2.2). Let $(X,\mathcal{B},\mu,T)$ be the product system where $X=X_{1}\times\cdots\times X_{k}$, $\mathcal{B}$ is the $\sigma$-algebra on $X$ generated by sets of the form $A_{1}\times\cdots\times A_{k}$ with $A_{i}\in\mathcal{B}_{i}$ for each $i=1,\ldots,k$, $\mu$ is the product measure on $\mathcal{B}$, and $T:X\rightarrow X$ is the product transformation defined by $T(x_{1},\ldots,x_{k})=(T_{1}(x_{1}),\ldots,T_{k}(x_{k}))$.
@@
3 A general form of Theorem 1 in $\mathbb{Z}$.
The goal of this section is to prove Theorem 3.2 which forms a general variant of Theorem 1 and has Theorem 1 as a rather special corollary. The proof of Theorem 3.2 is presented in Subsection 3.2 and is based on an amplification of the ergodic ideas utilized in the previous section. Subsection 3.2 is preambled by Subsection 3.1 where we collect some basic facts which will be needed in the proof of Theorem 3.2.
@@ 3.1 Ergodic sequences
It turns out that by working with sequences satisfying formula (3.1) and by introducing some additional amplifications of the ideas presented in the proof of Theorem 1, one can obtain new variants, such as Theorem 1
@@
    norm ergodic [^5] Although we will not use this result, we mention in passing that $(a_{n})_{n=1}^{\infty}$ is norm ergodic if and only if for any Hilbert space $\mathcal{H}$ and unitary operator $U:\mathcal{H}\to\mathcal{H}$ which has no invariant elements in $\mathcal{H}$, $\lim_{N\rightarrow\infty}\left\lVert\frac{1}{N}\sum_{n=1}^{N}U^{a_{n}}f\right\rVert_{\mathcal{H}}=0.$ (3.3) if for any ergodic invertible measure-preserving system $(X,\mathcal{B},\mu,T)$ and any $f\in L^{2}$,
@@
Note that if a sequence $(a_{n})_{n=1}^{\infty}$ in $\mathbb{N}$ is pointwise good and norm ergodic then it is pointwise ergodic. Also, using ergodic decomposition, one can verify that if $(a_{n})_{n=1}^{\infty}$ is pointwise ergodic, then it is simultaneously pointwise good and norm ergodic. Thus, when dealing with pointwise ergodic sequences, it is suitable to direct our attention to sequences which are pointwise good and norm ergodic.
@@
Theorem 3.2 ([2, Corollary 3]). A sequence $(a_{n})_{n=1}^{\infty}$ in $\mathbb{N}$ is norm ergodic if and only if for all $\lambda\in\mathbb{R}\setminus\mathbb{Z}$,
@@
Remark 3.3. It follows from Theorem 3.1 along with the Weyl criterion ([12, Chapter 1, Theorem 2.1] and [12, Chapter 5, Corollary 1.1]) that a sequence $(a_{n})_{n=1}^{\infty}$ in $\mathbb{N}$ is norm ergodic if and only if both of the following hold:
@@
Equivalently, $(a_{n})_{n=1}^{\infty}$ is norm ergodic if and only if equation (3.4) holds for two types of measure preserving systems: irrational rotations on the unit circle, and transformations of the form $x\mapsto x+1$ on $\mathbb{Z}/k\mathbb{Z}$ for $k\in\mathbb{N}$.
@@
Proof.
Imitating the proof of Theorem 5.12 in [7], we will first prove that $([g(n)]\gamma)_{n=1}^{\infty}$ is uniformly distributed mod 1 for $\gamma$ irrational. For all $(a,b)\in\mathbb{Z}^{2}\setminus\{(0,0)\}$ the sequence $(ag(n)\gamma+bg(n))_{n=1}^{\infty}=((a\gamma+b)g(n))_{n=1}^{\infty}$ is uniformly distributed mod 1, so $(g(n)\gamma,g(n))_{n=1}^{\infty}$ is uniformly distributed mod 1 ([12, Chapter 1, Theorem 6.3]). Then for all $h\in\mathbb{Z}\setminus\{0\}$ we have
$\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}e^{2\pi ih[g(n)]\gamma}$ $=\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}e^{2\pi ih(g(n)\gamma-\{g(n)\}\gamma)}$
@@
By [12, Chapter 5, Theorem 1.4], since $(g(n)/m)_{n=1}^{\infty}$ is uniformly distributed mod 1 for all integers $m\geq 2$ the sequence $([g(n)])_{n=1}^{\infty}$ is uniformly distributed in $\mathbb{Z}$. It follows from Remark 3.1 that $([g(n)])_{n=1}^{\infty}$ is norm ergodic. ∎
We will now describe a wide class of functions which, with the help of Theorem 3.1, provide a wide variety of norm ergodic sequences of the form $([g(n)])_{n=1}^{\infty}$.
@@
Theorem 3.5 ([10, Theorem 1.3]). Let $g\in\mathbf{U}$ be a subpolynomial function. Then the following two conditions are equivalent:
@@
    The sequence $(g(n))$ is uniformly distributed mod 1.

### v2

@@
Abstract. We apply the methods of ergodic theory to both simplify and significantly extend some classical results due to Stewart, Tijdeman, and Ruzsa. One of the notable features of our approach is the utilization of pointwise ergodic theory.
@@ 1 Introduction
$\overline{d}(S)=\limsup_{N\rightarrow\infty}\frac{|S\cap\{1,\ldots,N\}|}{N},$
@@
$\underline{d}(S)=\liminf_{N\rightarrow\infty}\frac{|S\cap\{1,\ldots,N\}|}{N}.$
If $\overline{d}(S)=\underline{d}(S)$, then the common value is denoted by $d(S)$ and is called the natural density of $S$. Following Rusza, we introduce three types of difference sets; note that these sets are symmetric subsets of $\mathbb{Z}$.
@@
Suppose that $S_{1},\ldots,S_{k}\subseteq\mathbb{N}$ and $\overline{d}(S_{i})>0$ for $1\leq i\leq k$. In [27], Stewart and Tijdeman answered in the affirmative a question of Erdős by showing that the set $C\vcentcolon=\bigcap_{i=1}^{k}\Delta_{3}(S_{i})$ has positive lower density. Moreover, they showed that $C$ is syndetic, meaning that there are integers $m_{1},\ldots,m_{\ell}$ such that
@@
Stewart and Tijdeman’s result was amplified by Rusza as follows.
Theorem 1.1 ([24, Theorem 1 and Theorem 2]). If $S_{1},\ldots,S_{k}\subseteq\mathbb{N}$ satisfy $\overline{d}(S_{i})>0$ for each $i=1,\dots,k$, then there exists $S\subseteq\mathbb{N}$ such that $d(S)\geq\prod_{i=1}^{k}\overline{d}(S_{i})$ and $\Delta_{1}(S)\subseteq D\vcentcolon=\bigcap_{i=1}^{k}\Delta_{2}(S_{i})$. Furthermore, there are $m_{1},\ldots,m_{\ell}\in\mathbb{Z}$, where
@@
The goal of this paper is to offer an ergodic approach to Theorem 1 which not only provides an alternative proof, but also leads to various generalizations which range from establishing versions of Theorem 1 in the setup of countably infinite amenable cancellative semigroups [^4] Amenable cancellative semigroups form a natural framework where notions of density can be defined. A precise definition of amenability is given in subsection 4.2. to variants of Theorem 1 which involve modified versions of the sets $\Delta_{1}(S)$ and $\Delta_{2}(S)$ such as
$\Delta_{1,c}(S)=\{n\in\mathbb{N}:S\cap(S-[n^{c}])\neq\emptyset\}\quad\text{ and }\quad\Delta_{2,c}(S)=\{n\in\mathbb{N}:\overline{d}(S\cap(S-[n^{c}]))>0\}$ (1.3)
with $c>0$, or, more generally,
$\Delta_{1,g}(S)=\{n\in\mathbb{N}:S\cap(S-[g(n)])\neq\emptyset\}\quad\text{ and}\quad\Delta_{2,g}(S)=\{n\in\mathbb{N}:\overline{d}(S\cap(S-[g(n)]))>0\}$ (1.4)
for certain functions $g:\mathbb{N}\rightarrow[1,\infty)$ (in formulas (1.3) and (1.4), []delimited-[][$\cdot$][ ⋅ ] denotes the integral part). An interesting new variant of Theorem 1 is provided by the following theorem, which deals with the function $g(n)=n^{c}$ and is a special case of the much more general Theorem 3.2, to be proved in subsection 3.2.
@@
Theorem 1.2. Let $c>0$, $c\not\in\mathbb{N}$. If $S_{1},\dots,S_{k}\subseteq\mathbb{N}$ satisfy $\overline{d}(S_{i})>0$ for each $i=1,\dots,k$, then there is $S\subseteq\mathbb{N}$ with $d(S)\geq\prod_{i=1}^{k}\overline{d}(S_{i})$ and $\Delta_{1,c}(S)\subseteq D_{c}\vcentcolon=\bigcap_{i=1}^{k}\Delta_{2,c}(S_{i})$. Furthermore, we have that $\overline{d}(\Delta_{1,c}(S))\geq\prod_{i=1}^{k}\overline{d}(S_{i})$. In addition, the set $D_{c}$ is thick, meaning that for all $n\in\mathbb{N}$ there is an $a\in\mathbb{N}$ such that $\{a,a+1,\dots,a+n\}\subseteq D_{c}$.
The reader may wonder why we require $c\notin\mathbb{N}$ in the statement of Theorem 1. The reason for this (to be revealed in subsection 3.1) is that for $c\notin\mathbb{N}$, the sequence $([n^{c}])_{n=1}^{\infty}$ is pointwise ergodic (see Definition 3.1), which in turn leads to the rather sharp inequality $d(S)\geq\prod_{i=1}^{k}\overline{d}(S_{i})$ appearing in the statement of Theorem 1. However, even when $c\in\mathbb{N}$, we still have the following nontrivial statement which naturally complements Theorem 1 and is a special case of Theorem 3.2.
Theorem 1.3. Let $c\in\mathbb{N}$. If $S_{1},\ldots,S_{k}\subseteq\mathbb{N}$ satisfy $\overline{d}(S_{i})>0$ for each $i=1,\ldots,k$, then there is $S\subseteq\mathbb{N}$ with $d(S)>0$, and $\Delta_{1,c}(S)\subseteq D_{c}\vcentcolon=\bigcap_{i=1}^{k}\Delta_{2,c}(S_{i})$. Furthermore, the set $D_{c}$ is syndetic.
Theorem 1 is weaker than Theorem 1 in the sense that it does not provide a bound for $d(S)$. But, unlike Theorem 1, Theorem 1 guarantees syndeticity of the set $D_{c}$.
Remark 1.4. An interesting feature of Theorem 1 is that for any $c>1$ with $c\not\in\mathbb{N}$ the set $D_{c}=\bigcap_{i=1}^{k}\Delta_{2,c}(S_{i})$ is thick. One can show that if $c\not\in\mathbb{N}$, unlike in Theorem 1, then the set $D_{c}$ may not be syndetic. Also when $c\in\mathbb{N}$ the syndetic set $D_{c}=\bigcap_{i=1}^{k}\Delta_{2,c}(S_{i})$ in Theorem 1 may not be thick. A detailed discussion regarding syndeticity and thickness of sets of the form $D_{g}=\bigcap_{i=1}^{n}\Delta_{2,g}(S_{i})$ for $g:\mathbb{N}\to[1,\infty)$, that appear in Theorems 3.2 and 3.2 (which contain Theorems 1 and 1 as rather special corollaries), is presented in subsection 3.3.
@@
As mentioned above, one of the major goals of this paper is to demonstrate that ergodic methods allow one to amplify and generalize Theorem 1 to the setup of countable cancellative amenable semigroups. Theorem 1, formulated below and proved in Section 4, generalizes Theorem 1 in more than one way. First, Theorem 1 extends the setup from $\mathbb{Z}$ to a general countably infinite amenable group $G$. Second, for the sets $S_{1},\dots S_{k}$, which appear in the formulation of Theorem 1, it is assumed $\overline{d}_{\mathbf{F}_{i}}(S_{i})>0$ for $i=1,\dots k$, where the Følner sequences $\mathbf{F}_{1},\dots,\mathbf{F}_{k}$ are arbitrarily chosen. Finally, the tempered Følner sequence $\mathbf{G}$, which appears in the formulation of Theorem 1, leads even in the case $G=\mathbb{Z}$ to a nontrivial generalization of Theorem 1. We will discuss Følner and tempered Følner sequences in detail in subsection 4.2.
@@
Theorem 1.5 (Theorem 4.2, subsection 4.2). Let $G$ be a countably infinite amenable group with left Følner sequences $\mathbf{G},\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$, where $\mathbf{G}$ is left tempered. Let $S_{1},\ldots,S_{k}\subseteq G$ such that $\overline{d}_{\mathbf{F}_{i}}(S_{i})>0$ for each $i=1,\ldots,k$. Then there is $S\subseteq G$ such that $d_{\mathbf{G}}(S)\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$ and $\Delta_{1}(S)\subseteq D=\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i})$. Furthermore, there are $m_{1},\ldots,m_{\ell}\in G$, where
@@
The structure of this paper is as follows. The goal of Section 2 is to provide a short ergodic proof of Theorem 1 and to set up some ideas that will be further developed and amplified in subsequent sections. In Section 3, we obtain Theorems 3.2 and 3.2 dealing with the difference sets defined in (1.4), which contain, respectively, Theorems 1 and 1 as instantaneous corollaries. In Section 4, we will prove Theorem 1 and its more general version, Theorem 4.3, which pertains to cancellative amenable semigroups. Finally, in Section 5, we give examples of tempered Følner sequences in various semigroups such as $(\mathbb{N},+)$, $(\mathbb{N},\cdot)$, the Heisenberg group, and locally finite groups.
Acknowledgments: The authors thank John Griesmer and Saúl Rodríguez Martín for helpful communications.
@@ 2 Ergodic proof of Ruzsa’s Theorem
The proof of Theorem 1 we present in this section will utilize two facts from ergodic theory. The first of them is an extension of the classical Poincaré recurrence theorem. [^5] Poincaré’s recurrence theorem states that if $(X,\mathcal{B},\mu)$ is a probability space, $T$ a $\mathcal{B}$-measurable map $X\to X$ such that $\mu(T^{-1}A)=\mu(A)$ for all $A\in\mathcal{B}$, and $B\in\mathcal{B}$ with $\mu(B)>0$, then there is $n\in\mathbb{N}$ such that $\mu(B\cap T^{-n}B)>0$.
Theorem 2.1 ([5, Theorem 1.2]). Let $(X,\mathcal{B},\mu,T)$ be a measure-preserving system [^6] A measure-preserving system is a quadruple $(X,\mathcal{B},\mu,T)$, where $(X,\mathcal{B},\mu)$ is a probability space and $T$ is a $\mathcal{B}$-measurable map $X\to X$ such that for any $A\in\mathcal{B}$, one has $\mu(T^{-1}A)=\mu(A)$. A measure-preserving system $(X,\mathcal{B},\mu,T)$ is invertible if the map $T$ is invertible. and let $B\in\mathcal{B}$ with $\mu(B)>0$. Then there is a set $P\subseteq\mathbb{N}$ with $d(P)\geq\mu(B)$ such that for any $n_{1},\ldots,n_{r}\in P$ we have $\mu\left(B\cap T^{-n_{1}}B\cap\cdots\cap T^{-n_{r}}B\right)>0$.
@@
Theorem 2.3 (cf. [6, Theorem 1.1]). Suppose that $S\subseteq\mathbb{N}$ with $\overline{d}(S)>0$. Then there is an invertible measure-preserving system $(X,\mathcal{B},\mu,T)$ and a set $B\in\mathcal{B}$ with $\mu(B)=\overline{d}(S)$ such that
@@
Proof of Theorem 1. By Theorem 2, for each $i=1,\ldots,k$ there exists an invertible measure-preserving system $(X_{i},\mathcal{B}_{i},\mu_{i},T_{i})$ and a set $B_{i}\in\mathcal{B}_{i}$ with $\mu_{i}(B_{i})=\overline{d}(S_{i})$ such that $S_{i}$, $B_{i}$, and $T_{i}$ satisfy formula (2.2). Let $(X,\mathcal{B},\mu,T)$ be the product system where $X=X_{1}\times\cdots\times X_{k}$, $\mathcal{B}$ is the $\sigma$-algebra on $X$ generated by sets of the form $A_{1}\times\cdots\times A_{k}$ with $A_{i}\in\mathcal{B}_{i}$ for each $i=1,\ldots,k$, $\mu$ is the product measure on $\mathcal{B}$, and $T:X\rightarrow X$ is the product transformation defined by $T(x_{1},\ldots,x_{k})=(T_{1}(x_{1}),\ldots,T_{k}(x_{k}))$.
@@
3 Two general variants of Theorem 1 for subsets of the integers
The goal of this section is to prove Theorems 3.2 and 3.2, each of which can be seen as a general variant of Theorem 1 and which have, correspondingly, Theorems 1 and 1 as rather special corollaries. The proofs of Theorems 3.2 and 3.2 are presented in subsection 3.2 and are based on an amplification of the ergodic ideas utilized in the previous section. Subsection 3.2 is preambled by subsection 3.1, where we collect some basic facts which will be needed in the proof of Theorems 3.2 and 3.2. Finally, in subsection 3.3, we collect some examples that illustrate the subtle features of Theorems 3.2 and 3.2 which pertain to the phenomena of thickness and syndeticity.
@@ 3.1 Ergodic sequences
It turns out that by working with sequences satisfying formula (3.1) and by introducing some additional amplifications of the ideas presented in the proof of Theorem 1, one can obtain new variants of Theorem 1, such as Theorem 1
@@
    norm ergodic [^7] Although we will not use this result, we mention in passing that $(a_{n})_{n=1}^{\infty}$ is norm ergodic if and only if for any Hilbert space $\mathcal{H}$ and unitary operator $U:\mathcal{H}\to\mathcal{H}$ which has no invariant elements in $\mathcal{H}$, $\lim_{N\rightarrow\infty}\left\lVert\frac{1}{N}\sum_{n=1}^{N}U^{a_{n}}f\right\rVert_{\mathcal{H}}=0.$ (3.3) if for every ergodic invertible measure-preserving system $(X,\mathcal{B},\mu,T)$ and any function $f\in L^{2}$,
@@
Note that if a sequence $(a_{n})_{n=1}^{\infty}$ in $\mathbb{N}$ is pointwise good and norm ergodic, then it is pointwise ergodic. Also, using ergodic decomposition, one can verify that if $(a_{n})_{n=1}^{\infty}$ is pointwise ergodic, then it is simultaneously pointwise good and norm ergodic. Thus, when dealing with pointwise ergodic sequences, it is suitable to direct our attention to pointwise good sequences and norm ergodic sequences.
@@
Theorem 3.2 ([4, Corollary 3]). A sequence $(a_{n})_{n=1}^{\infty}$ in $\mathbb{N}$ is norm ergodic if and only if for all $\lambda\in\mathbb{R}\setminus\mathbb{Z}$,
@@
Remark 3.3. It follows from Theorem 3.1 along with the Weyl criterion ([21, Chapter 1, Theorem 2.1] and [21, Chapter 5, Corollary 1.1]) that a sequence $(a_{n})_{n=1}^{\infty}$ in $\mathbb{N}$ is norm ergodic if and only if both of the following hold:
@@
Equivalently, $(a_{n})_{n=1}^{\infty}$ is norm ergodic if and only if formula (3.4) holds for two types of measure-preserving systems: irrational rotations on the unit circle, and transformations of the form $x\mapsto x+1$ on $\mathbb{Z}/k\mathbb{Z}$ for $k\in\mathbb{N}$.
@@
Proof. Imitating the proof of [9, Theorem 5.12], we will first prove that $([g(n)]\gamma)_{n=1}^{\infty}$ is uniformly distributed mod 1 when $\gamma$ is irrational. For all $(a,b)\in\mathbb{Z}^{2}\setminus\{(0,0)\}$ the sequence $(ag(n)\gamma+bg(n))_{n=1}^{\infty}=((a\gamma+b)g(n))_{n=1}^{\infty}$ is uniformly distributed mod 1, so $(g(n)\gamma,g(n))_{n=1}^{\infty}$ is uniformly distributed mod 1 ([21, Chapter 1, Theorem 6.3]). Define $f(x,y):=e^{2\pi ih(x-\{y\}\gamma)}$ and observe that, for all $h\in\mathbb{Z}\setminus\{0\}$, we have
$\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}e^{2\pi ih[g(n)]\gamma}$ $=\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}e^{2\pi ih(g(n)\gamma-\{g(n)\}\gamma)}=\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}f(g(n)\gamma,g(n)),$
@@
By the Weyl criterion ([21, Chapter 1, Theorem 2.1]) it therefore follows that $([g(n)]\gamma)_{n=1}^{\infty}$ is uniformly distributed mod 1.
By [21, Chapter 5, Theorem 1.4], since $(g(n)/m)_{n=1}^{\infty}$ is uniformly distributed mod 1 for all integers $m\geq 2$ the sequence $([g(n)])_{n=1}^{\infty}$ is uniformly distributed in $\mathbb{Z}$. It follows from Remark 3.1 that $([g(n)])_{n=1}^{\infty}$ is norm ergodic. $\square$
We will now describe a rather general class of functions which, with the help of Theorem 3.1, provide a wide variety of norm ergodic sequences of the form $([g(n)])_{n=1}^{\infty}$.
@@
Theorem 3.5 ([14, Theorem 1.3]). Let $g\in\mathbf{U}$ be a subpolynomial function. Then the following two conditions are equivalent:
@@
    The sequence $(g(n))_{n=1}^{\infty}$ is uniformly distributed mod 1.